OpenStudy (anonymous):

How do I solve the linear systems by substitution? y=x-3 y=-2x-6

OpenStudy (solomonzelman):

this is the system. \(\LARGE\color{blue}{ \sf y=x-3 }\) \(\LARGE\color{blue}{ \sf y=-2x-6 }\) way 1 \(\LARGE\color{blue}{ \sf y=\color{red} { x-3 }}\) substitute the red into the second equation \(\LARGE\color{blue}{ \sf \color{red} { x-3 }=-2x-6 }\) \(\LARGE\color{blue}{ \sf x-3 =-2x-6 }\) add 6 to both sides, \(\LARGE\color{blue}{ \sf x+3 =-2x }\) subtract x from both sides, \(\LARGE\color{blue}{ \sf 3 =-3x }\) divide both sides by -3 , \(\LARGE\color{blue}{ \sf ? =x }\) can you tell me the result ?

OpenStudy (solomonzelman):

way 2 is just to substitute from second equation to the first

OpenStudy (solomonzelman):

or substitute y instead of x.